OpenStudy (anonymous):

A container contains red, white, and yellow balls. Suppose that once a ball is drawn it is not replaced. What is the probability that red balls will be drawn in tries? (Express your answer as a fraction.)

OpenStudy (anonymous):

I think your Numlock key acted up

OpenStudy (anonymous):

A container contains 2 red, 3 white, and 5 yellow balls. Suppose that once a ball is drawn it is not replaced. What is the probability that 2 red balls will be drawn in 2 tries? (Express your answer as a fraction.)

OpenStudy (anonymous):

2 red out of 10 possibilities, then 1 out of 9. Mathematically, that's 2/10*1/9 = 1/45

OpenStudy (savvy):

(2/10)X(1/9) = 1/45